OpenStudy (anonymous):

What is 3 root of − 375 end root? A. − 3 3 root of 5 B. 5 3 root of − 9 end root C. 3 3 root of − 9 end root D. − 5 3 root of 3

OpenStudy (anonymous):

@thomas5267

OpenStudy (anonymous):

I think D

OpenStudy (anonymous):

@iGreen

OpenStudy (welshfella):

can you restated the question I don't get it

OpenStudy (anonymous):

A. \[-3 \sqrt[3]{5}\]

OpenStudy (welshfella):

do you mean the cube root of -375

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

B. 5 \[\sqrt[3]{-9}\]

OpenStudy (anonymous):

C. \[3 \sqrt[3]{-9}\]

OpenStudy (anonymous):

D. \[-5 \sqrt[3]{3}\]

OpenStudy (welshfella):

cube root of -375 = cube root(3 * -125) = -5 cube root 3

OpenStudy (welshfella):

because -5*-5*-5 = -125

OpenStudy (anonymous):

Ik but then you have to do this -125 x 3= -375

OpenStudy (welshfella):

yes - i did that

OpenStudy (welshfella):

i factored 375 into 3*125 because i knew 125 was a perfect cube

OpenStudy (anonymous):

oh sorry i didn't see it at first XD i'm still half asleep

OpenStudy (anonymous):

Thank you

OpenStudy (welshfella):

thats ok lol